OpenStudy (anonymous):

What is the solution of the equation? (x - 2)^3/4 = 8

OpenStudy (anonymous):

ok i see:)

OpenStudy (anonymous):

(x - 2)^3/4 = 8 Cube root both sides first (x-2)/4 = 2 Are you with me?

OpenStudy (anonymous):

yes:) i see

OpenStudy (anonymous):

Now you multiply each side by 4 (x-2) = 8 Add 2 to each side x= 10 That's about it :)

OpenStudy (anonymous):

wow that was easy. wish i could have done that on my own lol. Could you help me with one more? :)

OpenStudy (anonymous):

49. For the function f(x) = (3 - 4x)^2, find f^-1. Determine whether f^-1 is a function. This is my last question. :)